If 99 people have a mean income of Birr 400, how much is the mean income increased by the addition of a
woman with an income of Birr 1500? A. Birr 412 B. Birr 410 C. Birr 411 D.Birr413
1
Expert's answer
2020-06-08T18:52:56-0400

"The\; mean =\frac{99\times 400 +1500}{100}=411\\
Answer\;is\;C"